# StickerConv: Generating Multimodal Empathetic Responses from Scratch

### 20 Jan 2024 v1

@@ Abstract
Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research. In this paper, we introduce the Agent for StickerConv (Agent4SC), which uses collaborative agent interactions to realistically simulate human behavior with sticker usage, thereby enhancing multimodal empathetic communication. Building on this foundation, we develop a multimodal empathetic dialogue dataset, StickerConv, which includes 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios, specifically designs to augment the generation of empathetic responses in a multimodal context. To leverage the richness of this dataset, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation model, complemented by a comprehensive set of empathy evaluation metrics based on LLM. Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems. Our project page is available at https://neu-datamining.github.io/StickerConv.
StickerConv: Generating Multimodal Empathetic Responses from Scratch
@@ 1 Introduction
Figure 1: An example of multimodal conversation in the StickerConv. Both parties can utilize the stickers to express their emotions, which enhances interactivity and expression. Assistant can empathize with the user according to the conversation (green text).
Acknowledging stickers as a crucial non-textual modality and incorporating them into empathetic response generation offers novel perspectives and methodologies for empathetic dialogue tasks Zech and Rimé (2005); Rashkin et al. (2019). Specifically, integrating text and stickers paves the way for a new multimodal empathetic communication approach, targeting the generation of more diverse and high-quality empathetic responses.
A primary challenge in integrating stickers into empathetic response generation is developing a high-quality dataset to support this innovative multimodal communication. Traditional dataset construction methods, including expensive human crowdsourcing, privacy issues Xu et al. (2022), the difficulty in acquiring data from social software companies, and the variable quality of Internet-crawled data Foerderer (2023), face significant limitations.
@@
To address this, leveraging large language models (LLMs) for dataset construction is proposed. LLMs, with their extensive world knowledge and text processing capabilities (OpenAI, 2023; Touvron et al., 2023a, b), demonstrate near-human annotation abilities Wang et al. (2023d); Pangakis et al. (2023). However, LLMs have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking in proactivity Wu et al. (2023a); Yin et al. (2023), a critical aspect of empathy. Empathy necessitates understanding others’ emotions and the ability to actively express support and understanding Zech and Rimé (2005); Sharma et al. (2020a). To mitigate this, we introduce a multi-agent system based on LLMs, Agent for StickerConv (Agent4SC). This system, through inter-agent interactions, utilizes stickers to simulate human-like dialogue scenarios. It not only generates text responses but also strategically selects suitable stickers, thereby effectively enhancing empathy.
@@
Based on the superior capabilities and interactions of Agent4SC, we built a multimodal empathetic dataset, StickerConv, that comprise 12.9K dialogue sessions and 5.8K unique stickers. StickerConv boasts an average of 5.22 stickers per dialogue session. Figure 1 depicts an example of conversations in our dataset. To the best of out knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
@@
Although Agent4SC exhibits sound multimodal empathetic dialog response generation capability, it is limited by expensive inference costs, while the retrieval approach leads to poor performance in new scenarios. To further advance the research on multimodal empathetic dialog, based on StickerConv, we develop a end-to-end large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers, and has a significantly lower deployment and inference cost. Beyond the general ability to generate textual empathetic responses, PEGS receives multimodal inputs and autonomously generates stickers based on the emotional and contextual aspects of the dialogue at the appropriate moment. Furthermore, to simulate human communications on social media in the real world, our model supports interleaved multiple image and text inputs.
Due to potential misalignments between the modality quantities of predicted responses and those of the golden responses, reliance on a singular modal metric may lead to significant distortions in evaluation results. Empathy, being inherently subjective, presents challenges in its direct quantification. In this context, we introduce a novel multimodal empathetic response evaluation method. This method enables the integration of multimodal information, emphasizing three critical dimensions: empathy, consistency, and preference ranking. Leveraging the world knowledge and anthropomorphic capabilities of LLMs, our method offers robust support for evaluating multimodal empathetic responses.
@@
  • We introduce Agent for StickerConv (Agent4SC), a LLM-based multi-agent system, integrates stickers intelligently into empatheric dialogues. Agent4SC ensures the stickers are contextually consistent, varied, and empathetically aligned with human interaction. Leveraging the Agent4SC, we built a multimodal empatheric dialogue dataset, StickerConv, comprising 12.9K dialogue sessions and 5.8K unique stickers.
  • We design PErceive and Generate Stickers (PEGS), an empathetic dialogue response generation model that, within its end-to-end framework, intuitively incorporates stickers based on the emotional and contextual dynamics of a dialogue. PEGS adeptly processes multimodal inputs, generating empathetic textual responses and appropriately using stickers to enrich these replies.
  • We propose a multimodal empathetic response evaluation approach. This method integrates multimodal information and concentrates on three principal dimensions: empathy, consistency, and preference ranking.
@@ 2.1 Empathetic Response Generation
The purpose of empathetic response generation is to enable machines to understand and respond to human emotions and feelings. In this field, The creation of the EMPATHETICDIALOGUES dataset Rashkin et al. (2019) laid the foundation. Subsequently, the empathetic listener mixture model Lin et al. (2019), which emphasizes understanding and appropriately responding to different emotions, and studies focusing on emotion imitation and emotion cause words to enhance the effectiveness of empathetic dialogue Majumder et al. (2020) have achieved notable results. Additionally, the introduction of the EPITOME conceptual framework Sharma et al. (2020b), a classification system for empathetic response intents Welivita and Pu (2020), research on generating empathetic responses based on commensense knowledge Sabour et al. (2021), and the proposal of the multifactorial hierarchical framework CoMAE Zheng et al. (2021) have provided new perspectives and methods for the field.
Despite these advancements, the research in this domain is not without its limitations. Primarily, the field has been domainated by small-scale models, with limited application of large language models, such as ChatGPT OpenAI (2023), in the generation of empathetic responses Lee et al. (2022); Zhou et al. (2023b). Secondly, the focus has predominantly been on text-based empathetic responses, thus neglecting the rich potential of multimodal information, such as image. Emerging studies underscore the significance of such multimodal data, especially stickers, in alleviating stress, enhancing well-being, and fostering emotional resonance and social connectivity Kariko and Anasih (2019); Akram and Drabble (2022); Jiang et al. (2023). Finally, the subjectivity and multidimensionality of empathetic tasks make it challenging to quantitatively assess their effectiveness Fu et al. (2023); Lee et al. (2022), especially in the case of multimodal empathetic responses, posing a significant challenge in the field of empathetic response generation.
@@ 2.2 Large Multimodal Models
Multimodal Perception
Large language models (LLMs), such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains. Flamingo (Alayrac et al., 2022) exhibits promising zero-shot and few-shot multimodal understanding capability by adding a cross-attention layer to connect the frozen vision encoder with the LLM. BLIP (Li et al., 2023b; Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023) and LLaVA (Liu et al., 2023b, a) bridge the frozen vision encoder and the LLM through a small intermediate model.
Multimodal Generation
The birth and development of the Diffusion Models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Rombach et al., 2022) significantly contribute to the evolution of image generation. Koh et al. (2023) for the first time explore the mapping of the output of LLMs into the input space of the vision decoder, empowering LLMs with image generation capability. GILL (Koh et al., 2023), NExT-GPT (Wu et al., 2023b), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former. In contrast to previous works which favor realistic images, our target is to generate stickers, which are abstract and exhibit distinct emotional tendencies.
@@ 2.3 LLM-Based Agents
In the realm of artificial intelligence, Large Language Model (LLM)-based agents epitomize a cutting-edge paradigm, leveraging the robust computational prowess of LLMs. These agents are not merely confined to problem-solving and task execution; their capabilities extend to intricate reasoning and sophisticated interactions, as elucidated in recent studies Wang et al. (2023b); Li et al. (2023a). The versatility of LLM agents is further demonstrated through their adaptability in diverse domains such as software development Qian et al. (2023) and scientific research Boiko et al. (2023).
A remarkable aspect of LLM-based agents is their ability to emulate complex human social behaviors. This includes a spectrum of interactions ranging from social engagements Park et al. (2023); Tu et al. (2023), to simulating multifaceted scenarios like elections Argyle et al. (2022), debates Wang et al. (2023a), and consumer activities Wang et al. (2023c), Such capabilities underscore the potential of LLM agents in understanding and replicating human social dynamics.
The most transformative application of LLM-based agents, however, lies in their utility for generating training data for diverse AI models. Studies Wang et al. (2023d); Peng et al. (2023); Zhou et al. (2023a) have been instrumental in showcasing the efficacy of these agents in producing high-quality data, paralleling human annotators, yet at a substantially lower cost. This breakthrough has redefined the paradigm for creating training datasets, offering a more practical and cost-effective solution in the field of AI model development.
Figure 2: The Overview of Agent for StickerConv.
@@ 3 Agent for StickerConv
The limitations of Large Language Models (LLMs) and Language Model Multimodalities (LMMs), are twofold: (1) They struggle with capturing nuanced human emotions and providing empathetic responses. (2) While adept at responding to explicit human instructions, they lack the initiative to act beyond these directives.
This inherent disadvantage hinders their performance in tasks that require empathy, making them less effective in this area compared to others. To bridge this gap, we propose the Agent for StickerConv (Agent4SC), a novel LLM-based multi-agent. This system is designed to simulate human dialogue by leveraging the collaborative efforts of multiple modules, along with the use of stickers, to produce emotional and diverse empathetic responses.
The architecture of Agent4SC is depicted in Figure 2 and comprises two key agents. The User Agent, forming the core of the system, is an intricate assembly of several modules. It replicates the human processes of observation and reflection through the interplay between its Memory and Plan modules. This design effectively addresses the LLMs’ shortcomings in comprehending subtle human emotions. The insights gained from observation and reflection are articulated via the Action module. Consequently, the Agent is more adept at generating human-like emotional responses. The entire process is orchestrated by the Profile module, which imparts distinctiveness in reflection and behavior to different User Agents.
Furthermore, Agent4SC adopts a pioneering approach by treating stickers as a Tool of communication. Agents autonomously select appropriate stickers for use, mirroring human behavior in conversations. This function is facilitated through the integrated process of observation, reflection, and action. The User Agent ensures consistent performance and quality control with the support of the Manager Agent. Agent4SC is not just a static tool but a dynamic participant in conversations, capable of grasping the subtleties of human emotions and the appropriate moments for empathetic interaction.
@@ 3.1 Profile Module
The profile module serves as the foundation of the Agent4SC, shaping the personality and behavioral patterns of users in empathetic interactions. It comprises two components: Persona and Situation. The Persona section vividly portrays the user’s character traits, life background, and experiences, whereas the Situation part focuses on the user’s current circumstances, emotional state, and potential challenges they may encounter.
To develop rich and diversified user profiles, we initially manually crafted profiles with a variety of emotioinal distributions, subsequently expanding them to 2K unique profiles through the SELF-INSTRUCT Wang et al. (2023d) method by ChatGPT-3.5-turbo. Our constructed profile archive features a comprehensive emotional distribution, as illustrated in Figure 5. We believe that emotionally rich and content-diverse user profiles will enable the agent system to more authentically simulate human responses in empathetic dialogues.
@@ 3.2 Tool Module
To effectively simulate human behavior in using stickers for communication, we developed a Tool Module, converting the SER30K dataset into a functional tool. This dataset comprises 1,887 themes and 30,739 emotion-labeled stickers (Liu et al., 2022), such as happiness, anger, neutral, disgust, fear, sadness and surprise. In integrating it with the agent system, we encountered several challenges: (1) The dataset contains a significant number of stickers unsuitable for casual conversation or empathetic communication. (2) Stickers are annotated with a single emotional tag from a set of seven, lacking comprehensive content descriptions and a detailed emotional analysis. This limitation hinders the agent’s ability to appropriately utilize stickers based solely on these tags. (3) Humans selecting stickers in conversation typically do not choose from a database containing tens of thousands of sticker. Instead, they prefer to pick from a customized and smaller-scale sticker library.
To overcome these challenges, we design a series of sophisticated sticker processes to refashion sticker into an effective tool for accurate agent application:
Filter
Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we crafted a nuanced three-step Chain of Thought Wei et al. (2023) procedure to meticulously scrutinize each sticker in SER30K, as shown in Figure 9. This approach entailed the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses. This strategy not only affirmed selection accuracy but also mitigated model-induced randomness and hallucination discrepancies.
Annotation
For filtered stickers, a second iteration of the CoT process was executed using the same model. This phase involved responding to five pivotal queries, covering aspects like the sticker’s content, emotional conveyance, applicable scenarios, emotional tone, and potential satirical or humorous connotations (shown in Figure 9). These responses yielded comprehensive insights into each sticker, subsequently formatted into Q&A pairs, awaiting further processing.
Knowledge
Post-acquisition of intricate sticker details, the ChatGPT-3.5-Turbo model was engaged for an additional processing phase. Relying on the established Q&A pairs, this model adeptly deduced each sticker’s specific contextual suitability, culminating in the distillation of three fundamental knowledge: description, emotion, recommendation. These synthesized insights are instrumental in enriching agent’s understanding of sticker, and enable them to apply stickers correctly.
Split
In this phase of refining the SER30K sticker dataset for integration into our agent system, we focus on dividing the database into smaller, more manageable subsets. Each subset is transformed into a vector database, where the embeddings represent a unique blend of the original emotional labels of the stickers and the enriched emotional knowledge extracted in the previous steps.
After the above-mentioned processes, the original SER30K dataset has been transformed into a series of refined, smaller-scale vector databases. However, a crucial step is still required to make it a directly usable tool for agents. To address this, we have defined a specific module operation, termed "Retrieve". In this process, when an agent intends to use a sticker, it queries its accessible sub-vector database with its intended use and desired emotional expression of the sticker as search criteria. The system then retrieves and returns the top-K stickers with the highest similarity.

### 16 Feb 2024 v2

@@ Abstract
Stickers, while widely recognized for enhancing empathetic communication in online interactions, remain underexplored in current empathetic dialogue research, notably due to the challenge of a lack of comprehensive datasets. In this paper, we introduce the Agent for STICKERCONV (Agent4SC), which uses collaborative agent interactions to realistically simulate human behavior with sticker usage, thereby enhancing multimodal empathetic communication. Building on this foundation, we develop a multimodal empathetic dialogue dataset, STICKERCONV, comprising 12.9K dialogue sessions, 5.8K unique stickers, and 2K diverse conversational scenarios. This dataset serves as a benchmark for multimodal empathetic generation. To advance further, we propose PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation framework, complemented by a comprehensive set of empathy evaluation metrics based on LLM. Our experiments demonstrate PEGS’s effectiveness in generating contextually relevant and emotionally resonant multimodal empathetic responses, contributing to the advancement of more nuanced and engaging empathetic dialogue systems [^1] https://github.com/ZhangYiqun018/StickerConv.
STICKERCONV: Generating Multimodal Empathetic Responses
from Scratch
@@ 1 Introduction
Figure 1: An example of multimodal conversation in the STICKERCONV. Both parties can utilize the stickers to express their emotions, which enhances interactivity and expression. The assistant can empathize with the user according to the conversation (green text).
@@
Integrating stickers with textual communication and interspersing stickers within the dialogue can yield more varied and superior-quality empathetic replies. A primary challenge in integrating stickers into empathetic response generation is developing a high-quality dataset to support this innovative multimodal communication. To address this, we leverage large language model (LLMs) for dataset construction. LLMs, with their extensive world knowledge and text processing capabilities (OpenAI, 2023; Touvron et al., 2023a, b), demonstrate near-human annotation abilities Wang et al. (2023d); Pangakis et al. (2023). However, applying LLMs directly have limitations in empathetic tasks, excelling in responding to explicit human instructions but lacking proactivity, a critical aspect of empathy Wu et al. (2023); Yin et al. (2023). Empathy necessitates understanding others’ emotions and the ability to actively express support and understanding Zech and Rimé (2005); Sharma et al. (2020). To mitigate this, we introduce a multi-agent system based on LLMs, Agent for STICKERCONV (Agent4SC). This system, through inter-agent interactions, utilizes stickers to simulate human-like dialogue scenarios. It not only generates text responses but also strategically selects suitable stickers, thereby effectively enhancing empathy.
@@
Based on Agent4SC, we build a multimodal empathetic dataset, STICKERCONV, that comprises 12.9K dialogue sessions and 5.8K unique stickers. STICKERCONV boasts an average of 5.22 stickers per dialogue session, mirroring the sticker usage patterns observed in human communication. Figure 1 depicts an example of conversations in our dataset. To the best of our knowledge, this is the first multimodal empathetic dialogue dataset, with the particular utility of sticker as non-textual modal information to better facilitate empathy.
@@
Although Agent4SC effectively generates multimodal empathetic responses, it is limited by expensive inference costs and specific sticker databases. To further advance the research on multimodal empathetic dialogue, we develop an end-to-end multimodal empathetic response generation framework, PEGS, with the ability to PErceive and Generate Stickers. Beyond the general ability to generate textual empathetic responses, PEGS receives multimodal inputs and autonomously generates stickers based on the emotional and contextual aspects of the dialogue at the appropriate moment. Furthermore, to simulate human communications on social media in the real world, our model supports interleaved multiple image and text inputs.
Misalignments between the modal quantities of predicted and golden responses can distort evaluation outcomes, and empathy is difficult to quantify due to its subjective nature. To address this, we propose a novel method for evaluating multimodal empathetic responses, focusing on empathy, consistency, and ranking. Utilizing the extensive world knowledge and anthropomorphic abilities of LLMs, this approach provides solid support for assessing multimodal empathetic replies.
@@
  • We introduce an LLM-based multi-agent system, Agent for STICKERCONV (Agent4SC), which integrates stickers into empathetic dialogues, ensuring contextual consistency, variety, and empathy aligned with human interactions. Using Agent4SC, we create a multimodal empathetic dialogue dataset, STICKERCONV.
  • We design PErceive and Generate Stickers (PEGS), a multimodal empathetic response generation framework that intuitively incorporates stickers based on the emotional and contextual dynamics of the dialogue. PEGS adeptly processes multimodal inputs, generating empathetic textual responses and using stickers appropriately to enhance these responses.
  • We propose a method for assessing multimodal empathetic responses. It leverages LLM to evaluate the quality of these responses, with a specific focus on empathy, consistency, and ranking.
@@ 2.1 Empathetic Response Generation
Empathetic response generation focuses on enabling machines to understand and respond to human emotions. The foundational EMPATHETICDIALOGUES dataset Rashkin et al. (2019) and subsequent innovations like the empathetic listener mixture model Lin et al. (2019) have significantly advanced this area. Large Language Models (LLMs) like ChatGPT have been explored for empathetic response generation, although their application remains limited Lee et al. (2022); Zhou et al. (2023b). Challenges persist, especially in leveraging multimodal information for richer emotional engagement and in accurately evaluating empathetic responses due to their subjective nature. These obstacles highlight the ongoing need for research in effectively integrating LLMs and multimodal data into empathetic dialogue. Additionally, the subjective nature of empathy complicates its quantitative assessment Fu et al. (2023); Lee et al. (2022), posing a further obstacle to the field’s advancement.
@@ 2.2 Large Multimodal Models
LLMs, such as ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023a, b), demonstrate powerful capabilities in dialog interaction and instruction following, and recent researches have extended LLMs to multimodal domains. Flamingo (Alayrac et al., 2022) exhibits promising zero-shot and few-shot multimodal understanding capability by adding a cross-attention layer to connect the frozen vision encoder with the LLM. BLIP (Li et al., 2023b; Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023) and LLaVA (Liu et al., 2023b, a) bridge the frozen vision encoder and the LLM through a small intermediate model. Koh et al. (2023) for the first time explore the mapping of the output of LLMs into the input space of the vision decoder, empowering LLMs with image generation capability. GILL (Koh et al., 2023), MiniGPT-5 (Zheng et al., 2023) align LLMs to frozen vision decoders through an encoder-decoder transformer, while SEED (Ge et al., 2023) employs a learnable Q-Former. In contrast to previous works which favor realistic images, our target is to generate stickers, which are abstract and exhibit distinct emotional tendencies.
@@ 2.3 LLM-Based Agents
LLM-based agents mark a major leap in AI, leveraging their capabilities for tasks like reasoning and interaction, as shown in recent studies Wang et al. (2023b); Li et al. (2023a). They find uses across various domains, such as software engineering Qian et al. (2023) and scientific inquiry Boiko et al. (2023), highlighting their versatility. These agents can imitate complex human actions, partake in social interactions Park et al. (2023); Tu et al. (2023), and replicate intricate scenarios like elections Argyle et al. (2022), debates Wang et al. (2023a), and consumer patterns Wang et al. (2023c), illustrating their capacity to emulate human social dynamics. A notable innovation is their use in generating AI training data. Studies Wang et al. (2023d); Peng et al. (2023); Zhou et al. (2023a) highlight their efficiency and cost-effectiveness in producing high-quality training datasets, transforming AI model development.
Figure 2: The overview of Agent4SC. Memory and Plan modules enable the agent to mimic human observation and thought, overcoming LLMs’ inability to grasp nuanced emotions. The Action module supports generating insights with human-like emotional reactions. The Profile module gives each agent distinct reflections and actions. Furthermore, Agent4SC uses stickers as a Tool for more natural conversation, allowing the agent to choose stickers like humans. These modules streamline observation, reflection, and action, while the Manager Agent maintains performance and quality.
@@ 3 Agent for STICKERCONV
Confronted with the pivotal challenge of a lack of datasets for multimodal empathetic response tasks, we craft our own dataset utilizing large language models (LLMs). Nevertheless, LLMs experience difficulties in grasping nuanced human emotions and initiating actions beyond explicit directives. These limitations render LLMs and large multimodal models (LMMs) less proficient in the generation of multimodal empathetic responses. In response to these issues, we introduce Agent for STICKERCONV (Agent4SC), a multi-agent system predicated on LLM, devised to mimic human conversational patterns. Figure 2 presents the overview of Agent4SC. By integrating multiple modules and the strategic use of stickers, Agent4SC aims to generate emotional and varied empathetic responses, thereby overcoming the inherent deficiencies of LLMs in empathetic engagements.
@@ 3.1 Profile Module
The profile module, comprising Persona and Situation components, underpins its foundation by defining users’ personality traits and behavioral patterns for empathetic interactions. Persona outlines users’ character traits, backgrounds, and experiences, while Situation details their current circumstances and emotional states.
To enrich and diversify user profiles, we initially create profiles with varied emotional distributions, then expand to 2,000 unique profiles using SELF-INSTRUCT Wang et al. (2023d) method. This approach results in a profile archive with a broad emotional spectrum, as shown in Figure 5, aiming to enhance the system’s ability to simulate human-like responses in empathetic dialogues.
@@ 3.2 Tool Module
To adapt the SER30K dataset, with its 1,887 themes and 30,739 emotion-tagged stickers (Liu et al., 2022), for human-like sticker use in dialogues, we transform it into a tool. Integrating it into the agent system faced several hurdles: the inappropriateness of many stickers for empathetic dialogues, the lack of detailed content or emotional analysis in the single-emotion labeled stickers, and the contrast between the extensive SER30K collection and the limited, personalized collections humans use.
To make stickers an effective tool, we refine the SER30K through filtering, annotating, and knowledge extraction (detailed in Appendix B.1). Each sticker is enriched with three pieces of information: emotion, description, and recommendation. Subsequently, SER30K is segmented into smaller, emotionally balanced vector databases. A critical action, "Retrieve", is then designed for agents to efficiently use these databases by searching for stickers based on the communication context and desired emotion, with the system suggesting the top-K closest matches for selection.
3.3 Memory Module
In empathetic interactions, response quality hinges on communication coherence and personality trait consistency. Agents must learn and give feedback by observing during interactions and ensure long-term personality consistency. Following the Generative Agent Park et al. (2023), we use short-term updates of core traits from profiles and observations for immediate memory. For long-term memory, continuous interactions and self-reflection help maintain consistency.